If 10 grams of salt are present in 500 mL of saline solution, what is the percentage of salt in the solution?
For some reason, I came up with %=part/whole*100 as a possible way to solve the problem, but I could be wrong and am lost.
Thanks,

Answer by stanbon(75887)

If 10 grams of salt are present in 500 mL of saline solution, what is the percentage of salt in the solution?
----
10g/500mL = (1/50)g/mL = 0.005g/mL = 0.5%
===========================================
Cheers,
Stan H.